# Sample-efficient non-Gaussian noise reduction in gravitational wave data via learnable wavelets

### v1

@@ Abstract
Flexible machine learning (ML) architectures such as CNNs and transformer-based models have been widely applied to the identification of non-Gaussian noise in GW data Koloniari et al. (2025); Zhao et al. (2025); Yamamoto et al. (2022); Zevin and others (2024). While these approaches are highly expressive - i.e., they can learn and model complex, nonlinear relationships and hierarchical features in the data - they typically require large training datasets, particularly when operating on long time-domain segments (tens of seconds or more). This data demand can be a limiting factor in regimes where representative examples of specific glitch morphologies are scarce.
Empirically, a class of glitches that contaminate searches for high-mass binary black hole mergers exhibits localized, wavelet-like time-frequency structure. Motivated by this observation, we introduce an explicit inductive bias in our neural network method by modeling glitches using a parameterized wavelet family: the Morlet wavelets ($\psi(t)=e^{-t^{2}/2\sigma^{2}}e^{i\omega_{0}t}$), as an approximate generative model. Rather than learning arbitrary features from raw data, we employ a lightweight multi-layer perceptron (MLP) to infer physically interpretable wavelet parameters, such as central frequency and temporal duration. This approach trades architectural flexibility for data efficiency and interpretability, while remaining sufficiently expressive to capture the dominant glitch morphologies relevant to high-mass searches. We apply this framework to model and characterize glitches in data from the LIGO–Virgo–KAGRA O3 observing run.
@@
In this work, we introduce WaveletNet, designed to more generally characterize nonlocal strain data. WaveletNet incorporates learnable Morlet wavelets to construct glitch-like templates and performs matched filtering over an extended region surrounding candidate triggers. By explicitly modeling the time-frequency morphology of instrumental glitches, the wavelet-based approach provides a strong inductive bias that is much more sample efficient than generic CNNs when applied to nonlocal data. When integrated modularly into the IAS-HM search pipeline, WaveletNet yields improvements of up to $15\%$ in sensitive spacetime volume, demonstrating its effectiveness as a complementary ranking component rather than a replacement for existing pipelines.
We apply WaveletNet within the IAS-HM search framework, building on the TIER methodology introduced in Ref. Wadekar and others (2025). Instrumental glitches often “come with friends,” meaning they cluster in time and occur in elevated noise environments. While this behavior has previously been used only in a limited, manual way to reduce contamination in the GWTC-4 catalog (see Fig. 9 of Ref. Abac and others (2025)), the current IAS-HM pipeline evaluates candidate significance using only a narrow $\pm 0.1\,\mathrm{s}$ window around the trigger. As a result, informative structure in the surrounding strain data is ignored. WaveletNet addresses this limitation by producing an external score from nonlocal data that can be modularly combined with the existing ranking statistic.
@@ I.1 Motivation for an Inductive-Bias-Driven Architecture
Widely used deep learning approaches (such as CNNs) treat this as a representation learning problem, where the model must learn both the signal structure and the decision boundary from raw data. While flexible, this typically requires large models and datasets, and offers limited physical interpretability. In contrast, classical signal processing provides established tools, such as matched filtering, that are optimal for improving SNRs when both the signal shape and noise distribution is partially or entirely known. Our approach combines these ideas by building signal-processing structure directly into the early stages of the pipeline.
@@ II.1 Learnable Wavelet Matched Filtering Module
$\psi(t)=\pi^{\mathrm{-1/4}}\cdot e^{i2\pi ft}\cdot e^{-t^{2}/2}\,.$ (1)
@@
An analytical phase correction is applied to maximize the matched filter output. The required correction $\Delta\phi(t)$ is
@@
[…]
@@ II.2 MLP Classification Head
In this work, the ML component is designed to augment the traditional matched filtering search pipeline rather than replace it. The glitches used for training are noise transients that produce triggers in the existing template banks and therefore span a range of morphologies, particularly in their characteristic frequency content. Capturing this diversity with a single wavelet model is challenging. We thus train separate instances of the models for each template bank. Since the banks are approximately partitioned by total mass, and hence by characteristic signal frequency, glitches associated with a given bank tend to exhibit similar time-frequency structure. For each bank, the model is trained exclusively on glitches identified by matched filtering with templates from that bank. During inference, for a given candidate trigger, the external probability $p^{\mathrm{ext}}$ is computed using the model corresponding to the bank of the triggering template, ensuring consistency between the candidate’s frequency content and the learned glitch representation.
@@ III Performance on a toy dataset
In Fig. 4 we show the validation performance of the two models as a function of training dataset size and training epoch. The wavelet-based model achieves higher validation accuracy with significantly fewer training samples and exhibits stable convergence during training. In contrast, the CNN shows slower improvement with dataset size and substantial instability in validation loss, reflecting its higher data requirements and sensitivity to overfitting. These trends qualitatively demonstrate how incorporating problem-specific structure can improve data efficiency and training robustness, even in simplified settings.
Figure 4: Comparison of a wavelet-based neural network vs. a CNN on a toy-model dataset (see section III for details). This illustrates the effect of the inductive bias used in WaveletNet on data efficiency and training stability. Top panel: Validation accuracy as a function of training dataset size for a wavelet-based model and a CNN. The wavelet-based model achieves higher accuracy with fewer training samples. Bottom panel: Validation loss versus training epoch for the same models, showing stable convergence for the wavelet-based model and significantly larger fluctuations for the CNN. These results qualitatively demonstrate the advantage of incorporating signal-processing structure in limited-data regimes.
@@ IV Including environmental information in the search ranking statistic
We apply ML methods to address this issue and learn the distribution of non-Gaussian noise from the detector data and down-weight non-Gaussian transients Biswas and others (2013); Zevin and others (2017); Cuoco et al. (2025); Fernandes et al. (2025); Sharma et al. (2022); Saleem and others (2024); Reissel et al. (2025); Alvarez-Lopez et al. (2024); Lopez et al. (2025); Malz and Veitch (2025); Legin et al. (2025); Shah et al. (2023); Marx and others (2025); Schäfer and others (2023); McLeod et al. (2025); Kapadia et al. (2017); Koloniari et al. (2025); Bini et al. (2023); Essick et al. (2020); Nagarajan and Messenger (2025); Chatterjee et al. (2024); Chatterjee and Jani (2024, 2025); McIsaac and Harry (2022). As shown schematically in Fig. 5, the resulting ML score can be modularly combined with the traditional pipeline statistic to further calibrate the ranking score:
@@ V Training Data
In this section, we describe the input data streams used for training WaveletNet and obtaining $p^{\mathrm{ext}}_{i}$. WaveletNet integrates two inputs: a whitened strain time series from the nonlocal region surrounding each candidate, and a vector of statistical features summarizing the candidate’s properties. Together, these inputs provide complementary contextual and summary information. The following sections describe these inputs in detail. Separate models are trained for each of the 17 binary black hole template banks, whose construction and bank boundaries are described in Section II of Wadekar et al. (2024b).
@@ V.1 Whitened Time Series
V.2 Nonlocal Statistical Features of the Candidate
@@
We present below the statistical features computed from the nonlocal region of the data. These features were developed through empirical, trial and error exploration, with the aim of capturing characteristics of the surrounding data that may indicate elevated noise levels or nonstationarity. Similar summary statistics are also used in the random forest models described in Ref. Wadekar and others (2025). We note that the development of more informative summary statistics, or the use of ML based compression methods for the environmental data, may further improve classification performance.
@@
Figure 6: The two classes of inputs to WaveletNet are the SNR responses obtained from matched filtering with the learned wavelets, and the additional summary statistics provided by the search pipeline (see Fig. 1). Here, we illustrate the relative importance of these two input classes by quantifying their effect on the improvement in IFAR for injections in the data. The left panel shows the performance obtained using only the wavelet matched filtering outputs. The center panel shows the IFAR improvement obtained when using only the summary statistics from the search pipeline. The right panel shows the results for the full WaveletNet model, which combines both the wavelet-based features and the pipeline summary statistics. We find that the wavelet-based features provide a stronger contribution to the IFAR improvement than the pipeline summary statistics alone.
@@ VI Search pipeline Sensitivity Improvement
Our computation of the volume-time, $\overline{VT}$, employs a weighted Monte Carlo integral over the injections recovered by the pipeline with $\mathrm{IFAR}>1$ yr, following Tiwari (2018):
@@ VII Discussion
A key distinction between WaveletNet and TIER lies in how nonlocal noise is modeled. The TIER method relies on matched filtering with template banks originally constructed to identify compact binary coalescence signals. While this approach is effective at capturing certain forms of excess power, these templates are not optimized for instrumental glitches, which often exhibit time-frequency morphologies that differ substantially from those of astrophysical signals. In contrast, WaveletNet learns a set of data-driven wavelet templates directly from environmental data, enabling it to adapt to the characteristic time-frequency structure of glitches. This design choice allows the model to assess the noise environment surrounding a candidate in a manner more closely aligned with the underlying glitch phenomenology.
@@
In terms of detection efficiency, we find that the improvement in sensitive spacetime volume $VT$ achieved on O3 data is comparable to that obtained with the original TIER approach. While the gains are not substantially larger in this dataset, the wavelet-based method is complementary rather than redundant. Because WaveletNet learns glitch morphologies directly from the data, it is naturally suited to adapting to changes in detector behavior. We therefore expect this approach to be particularly valuable in future observing runs, such as O4 and beyond, where the population and morphology of glitches may differ from those seen in O3. As long as glitches admit a reasonably sparse wavelet representation, the method can retrain and recalibrate without relying on fixed signal-based template banks.
@@
We also explored several alternative ML approaches based on CNNs, including architectures augmented with attention mechanisms. These models offer greater flexibility and, in principle, can learn a broader class of time-frequency patterns than wavelet-based models. However, in practice, we find that CNN-based approaches require significantly more training data to achieve stable performance and are more prone to overfitting in the limited data regime relevant for environmental glitch characterization. While CNNs may ultimately outperform wavelet-based methods given sufficiently large and diverse training sets, WaveletNet provides a more data-efficient solution for the problem considered here. A detailed discussion of these alternative models and their performance is provided in Appendix A.
@@ VIII Conclusions
Search pipelines for GWs typically assess the origin of a candidate using data localized to the immediate vicinity of the event. However, information contained in the broader, nonlocal data environment can provide valuable additional context for distinguishing astrophysical signals from instrumental noise. In this work, we have developed ML models that analyze the nonlocal environment surrounding a candidate by matched filtering learnable wavelets against the strain data, complemented by sparse summary statistical features.
The scalar output of these models can be straightforwardly incorporated into existing ranking statistics used by search pipelines, enabling re-ranking of candidates without disrupting established detection workflows. Applying this approach to the IAS-HM search, we observe improvements of up to $\sim 15\%$ in the sensitive spacetime volume, as shown in Fig. 7, demonstrating enhanced detection efficiency at fixed false-alarm rate.
Importantly, the proposed framework is general and can be integrated with any GW search pipeline that produces candidate triggers and associated environmental data. Because the wavelet templates are learned directly from the data, the method is well suited to adapting to changes in detector noise characteristics. Future work will extend this analysis to O4 and subsequent observing runs, where evolving glitch populations and increased detector sensitivity are expected to further benefit from data-driven adaptive noise characterization.
@@ A.1 Convolutional Neural Network (CNN)
We experimented with a brute force method of directly training CNNs on nonlocal intervals with the local region of the candidate removed. We also tried including an attention transformer layer to focus on the glitches. We elaborate the same in the following sections.
@@ A.1.1 Training on spectrograms of the nonlocal region
We generated spectrograms from the time-series data using the specgram function from the matplotlib library. The spectrograms were computed with a sampling frequency Fs = 256, an FFT window size NFFT = 64, and an overlap of noverlap = 32 between successive windows. The spectrogram was divided into the left and right region of the candidate and passed separately to the CNN.

### v2

@@ Abstract
Flexible machine learning (ML) architectures such as CNNs and transformer-based models have been widely applied to the identification of non-Gaussian noise in GW data, see e.g., Koloniari et al. (2025); Zhao et al. (2025); Yamamoto et al. (2022); Zevin and others (2024); Gebhard et al. (2019); Li et al. (2017); Biswas and others (2013); Zevin and others (2017); Cuoco et al. (2025); Fernandes et al. (2025); Sharma et al. (2022); Saleem and others (2024); Reissel et al. (2025); Alvarez-Lopez et al. (2024); Lopez et al. (2025); Malz and Veitch (2025); Legin et al. (2025); Shah et al. (2023); Marx and others (2025); Schäfer and others (2023); McLeod et al. (2025); Kapadia et al. (2017); Bini et al. (2023); Essick et al. (2020); Nagarajan and Messenger (2025); Chatterjee et al. (2024); Chatterjee and Jani (2024, 2025); McIsaac and Harry (2022). While these approaches are highly expressive - i.e., they can learn and model complex, nonlinear relationships and hierarchical features in the data - they typically require large training datasets, particularly when operating on long time-domain segments (tens of seconds or more). This data demand can be a limiting factor in regimes where representative examples of specific glitch morphologies are scarce.
Empirically, a class of glitches that contaminate searches for high-mass binary black hole mergers exhibits localized, wavelet-like time-frequency structure. Motivated by this observation, we introduce an explicit inductive bias in our neural network method by modeling glitches using a parameterized wavelet family: the Morlet wavelets ($\psi(t)=e^{-t^{2}/2\sigma^{2}}e^{i\omega_{0}t}$), as an approximate generative model. Rather than learning arbitrary features from raw data, we employ a lightweight multi-layer perceptron (MLP) to infer physically interpretable wavelet parameters, such as central frequency and temporal duration. This approach trades architectural flexibility for data efficiency and interpretability, while remaining sufficiently expressive to capture the dominant glitch morphologies relevant to high-mass searches.
@@
In this work, we design a network, WaveletNet, to characterize noisy strain data in the vicinity of gravitational-wave candidates. WaveletNet incorporates learnable Morlet wavelets to construct glitch-like templates and performs matched filtering over an extended region surrounding candidate triggers. By explicitly modeling the time–frequency morphology of instrumental glitches, the wavelet-based approach provides a strong inductive bias that is much more sample efficient than generic CNNs when applied to non-local data. We apply our framework to model and characterize glitches in data from the LIGO–Virgo–KAGRA O3 observing run. It is also worth noting that wavelet-based approaches have been successfully applied in other areas of astrophysics and cosmology to achieve robust inference and improved data efficiency compared to generic deep neural networks (e.g., Ref. Cheng et al. (2020)).
When integrated modularly into the IAS-HM search pipeline, WaveletNet yields improvements of up to $15\%$ in sensitive spacetime volume, demonstrating its effectiveness as a complementary ranking component rather than a replacement for existing pipelines. This builds on the TIER methodology introduced in Ref. Wadekar and others (2025). Instrumental glitches often “come with friends,” meaning they cluster in time and occur in elevated noise environments [^1] Note that this behavior has previously been used only in a limited, manual way to reduce contamination in the GWTC-4 catalog (see Fig. 9 of Ref. Abac and others (2025)).. Typical current search pipelines evaluate candidate significance using only a narrow ($\sim\pm 0.1\,\mathrm{s}$) window around the trigger. As a result, informative structure in the surrounding strain data is ignored. WaveletNet addresses this limitation by producing an external score from nonlocal data that can be modularly combined with the existing ranking statistic.
@@ I.1 Motivation for an Inductive-Bias-Driven Architecture
Widely used deep learning approaches (such as CNNs) treat this as a representation learning problem, where the model must learn both the signal structure and the decision boundary from raw data. While flexible, this typically requires large models and datasets, and offers limited physical interpretability. In contrast, classical signal processing provides established tools, such as matched filtering, that are optimal for improving SNRs when both the signal shape and noise distribution is partially or entirely known. Our approach combines these ideas by using flexible signal-processing structure into the matched-filtering step of the pipeline.
@@ II.1 Learnable Wavelet Matched Filtering Module
$\psi(t)=\pi^{\mathrm{-1/4}}\cdot e^{i2\pi ft}\cdot e^{-t^{2}/2}\,$ (1)
@@
To maximize the matched-filter response, we apply an analytical phase maximization. This is equivalent to projecting the complex response onto its optimal phase, yielding the magnitude of the output,
@@
This operation maximizes the response over the relative phase between the data and the wavelet and is equivalent to an explicit phase rotation that aligns the complex output along the real axis.
@@ II.2 MLP Classification Head
In this work, the ML component is designed to augment the traditional matched filtering search pipeline, rather than replace it. The glitches used for training are noise transients that produce triggers in the existing GW template banks and therefore span a range of morphologies, particularly in their characteristic frequency content. Capturing this diversity with a single wavelet model is challenging. We thus train separate instances of the models for triggers from different GW template banks (the banks are approximately partitioned by total mass, see Wadekar et al. (2023) for details). The different banks correspond to waveforms with characteristic GW signal frequency, hence glitches associated with a given bank tend to exhibit similar time-frequency structure. Therefore, for each bank, the model is trained exclusively on glitches identified by matched filtering with templates from that bank. During inference, the probability of a given candidate trigger being a glitch is computed using the model corresponding to the bank of the triggering template.
@@ III Performance on a toy dataset
In Fig. 4 we show the validation performance of the two models as a function of training dataset size and training epoch. The wavelet-based model achieves higher validation accuracy with significantly fewer training samples and exhibits stable convergence during training. In contrast, the CNN shows slower improvement with dataset size and substantial instability in validation loss, reflecting its higher data requirements and sensitivity to overfitting. These trends qualitatively demonstrate how incorporating problem-specific structure can improve data efficiency and training robustness, even in simplified settings. Note that for very low training dataset size, there can be significant stochasticity in training the models, this is likely the reason for change in relative behavior for the in the left panel of Fig. 4.
More broadly, synergies between wavelet-based representations and flexible deep learning architectures have been explored in the machine learning literature, demonstrating that combining explicit time–frequency structure with learnable models can improve robustness and generalization in limited-data regimes Oyallon et al. (2019); Michau et al. (2022); Khemani et al. (2022); Bruna and Mallat (2012).
Figure 4: Comparison of a wavelet-based neural network and a CNN on a toy-model dataset (see Section III for details). This figure illustrates the effect of the inductive bias used in WaveletNet on data efficiency and training stability. Top panel: Validation accuracy as a function of training dataset size for a wavelet-based model and a CNN, showing that the wavelet-based model achieves higher accuracy with fewer training samples. Bottom panel: Validation loss as a function of training epoch for the same models, obtained using a fixed training set of 400400 examples per class. The wavelet-based model exhibits stable convergence, while the CNN shows significantly larger fluctuations. These results qualitatively demonstrate the advantage of incorporating signal-processing structure in limited-data regimes.
@@ IV Including environmental information in the search ranking statistic
We use our ML method to address this issue; we learn the distribution of non-Gaussian noise from the detector data and down-weight non-Gaussian transients. As shown schematically in Fig. 5, the resulting ML score can be modularly combined with the traditional pipeline statistic to further calibrate the ranking score:
@@ V Training Data
In this section, we describe the input data streams used for training WaveletNet and obtaining $p^{\mathrm{ext}}_{i}$. WaveletNet integrates two inputs: a whitened strain time series from the nonlocal region surrounding each candidate, and a vector of statistical features summarizing the candidate’s properties. Together, these inputs provide complementary contextual and summary information. The following sections describe these inputs in detail. Separate models are trained for each of the 17 binary black hole template banks, whose construction and bank boundaries are described in Section II of Wadekar et al. (2024b). For each bank, the models are trained using approximately 15,00015{,}000 injection examples and 23,00023{,}000 glitch examples.
@@ V.1 Whitened Time Series
V.2 Additional nonlocal statistical features of the GW candidates
@@
We present below the statistical features computed from the nonlocal region of the data. These features were developed through empirical, trial and error exploration, with the aim of capturing characteristics of the surrounding data that may indicate elevated noise levels or nonstationarity. Similar summary statistics are also used in the random forest models described in Ref. Wadekar and others (2025). We note that the development of more informative summary statistics, or the use of ML based compression methods for the environmental data, may further improve our ML classifier performance.
@@
Figure 6: The two classes of inputs to WaveletNet are the SNR responses obtained from matched filtering with the learned wavelets, and the additional summary statistics provided by the search pipeline (see Fig. 1). Here, we illustrate the relative importance of these two input classes by quantifying their effect on the improvement in IFAR for injections in the data. The left panel shows the performance obtained using only the wavelet matched filtering outputs. The center panel shows the IFAR improvement obtained when using only the summary statistics from the search pipeline (see section V.2). The right panel shows the results for the full WaveletNet model, which combines both the wavelet-based features and the pipeline summary statistics. We find that the wavelet-based features provide a stronger contribution to the IFAR improvement than the pipeline summary statistics alone.
@@ VI Search pipeline Sensitivity Improvement
Our computation of the volume-time, $\overline{VT}$, employs a more accurate weighted Monte Carlo integral over the injections recovered by the pipeline with $\mathrm{IFAR}>1$ yr, following Tiwari (2018):
@@ VII Discussion
A key distinction between WaveletNet and TIER lies in how nonlocal noise is modeled. The TIER method relies on matched filtering with template banks (originally constructed to identify compact binary coalescence signals). While this approach is effective at capturing certain forms of excess power, these templates are not optimized for instrumental glitches, which often exhibit time-frequency morphologies that can differ substantially from those of astrophysical signals. In contrast, WaveletNet learns a set of data-driven wavelet templates directly from environmental data, enabling it to adapt to the characteristic time-frequency structure of glitches. This design choice allows the model to assess the noise environment surrounding a candidate in a manner more closely aligned with the underlying glitch phenomenology.
@@
In terms of detection efficiency, we find that the improvement in sensitive spacetime volume $VT$ achieved on O3 data is comparable to that obtained with the original TIER approach. While the gains are not substantially larger in this dataset, the wavelet-based method is complementary rather than redundant. Because WaveletNet learns glitch morphologies directly from the data, it is naturally suited to adapting to changes in detector behavior. We therefore expect this approach to be particularly valuable in future observing runs, such as O4 and beyond, where the population and morphology of glitches may differ from those seen in O3. As long as glitches admit a reasonably sparse wavelet representation, the method can retrain and recalibrate without relying on fixed signal-based template banks. The wavelet parameters corresponding to glitches which are learnt by WaveletNet can be used as priors in parameter estimation and data cleaning methods such as BayesWave Cornish and Littenberg (2015).
@@
We also explored several alternative ML approaches based on CNNs, including architectures augmented with attention mechanisms. These models offer greater flexibility and, in principle, can learn a broader class of time–frequency patterns than wavelet-based models. However, in practice, we find that CNN-based approaches require significantly more training data to achieve stable performance and are more prone to overfitting in the limited-data regime relevant for environmental glitch characterization. For a fair comparison, the CNNs were trained using the same dataset sizes as WaveletNet, namely 15,00015{,}000 injection examples and 23,00023{,}000 glitch examples per bank. While CNNs may ultimately outperform wavelet-based methods given sufficiently large and diverse training sets, WaveletNet provides a more data-efficient solution for the problem considered here. A detailed discussion of these alternative models and their performance is provided in Appendix A.
@@ VIII Conclusions
A key takeaway of this work is that effective nonlocal noise characterization does not require highly complex model architectures. Instead of relying on large, data-hungry CNNs, we introduce a lightweight wavelet-based model that incorporates prior knowledge about glitch structure and can be trained with substantially fewer examples. This simpler architecture is easier to train, more stable in limited-data regimes, and still captures the features needed to improve candidate ranking.
Gravitational-wave search pipelines typically assess the origin of a candidate using data localized to the immediate vicinity of the event. However, information contained in the broader, nonlocal data environment can provide valuable additional context for distinguishing astrophysical signals from instrumental noise. In this work, we develop machine-learning models that analyze the nonlocal environment surrounding a candidate by matched filtering learnable wavelets against the strain data, complemented by sparse summary statistical features.
The scalar output of these models can be straightforwardly incorporated into existing ranking statistics used by search pipelines, enabling re-ranking of candidates without disrupting established detection workflows. Applying this approach to the IAS-HM search, we observe improvements of up to $\sim 15\%$ in sensitive spacetime volume, as shown in Fig. 7, demonstrating enhanced detection efficiency at fixed false-alarm rate.
Importantly, the proposed framework is general and can be integrated with any gravitational-wave search pipeline that produces candidate triggers and associated environmental data. Because the wavelet templates are learned directly from the data, the method is well suited to adapting to changes in detector noise characteristics. Future work will extend this analysis to O4 and subsequent observing runs, where evolving glitch populations and increased detector sensitivity are expected to further benefit from data-driven, adaptive noise characterization.
@@ A.1 Convolutional Neural Network (CNN)
We experimented with a brute force method of directly training CNNs on the GW strain data examples mentioned in section V. We also tried including an attention transformer layer to focus on the glitches. We elaborate the same in the following sections.
@@ A.1.1 Training on spectrograms of the nonlocal region
We generated spectrograms from the time-series data using the specgram function from the matplotlib library. The spectrograms were computed with a sampling frequency Fs = 256, an FFT window size NFFT = 64, and an overlap of noverlap = 32 between successive windows. The spectrogram was divided into the left and right region of the candidate and passed separately to the CNN (see Fig. 5).